ILoveXXXTENTACION:

Can u solve this for me, and just solve it not give me steps and just confuse me.

ILoveXXXTENTACION:

Tinkerbell:

2 x -7 = -14

Tinkerbell:

-14 x -7 = 98

Tinkerbell:

Figure the rest

Tinkerbell:

98 x -7 =? @ILoveXXXTENTACION

jhonyy9:

so the ratio is -7 how you get the 6th term ?

bubblegirl1234:

who like nba youngboy

ILoveXXXTENTACION:

yea, i like his song, "no smoke"